OpenStudy (animelover98):

Which of the following is a solution of x2 + 2x + 8? I know it is supposed to contain an Imaginary number, but I am having trouble getting through it. Can someone help me? Thank you!

OpenStudy (anonymous):

Do you know the quadratic formula?

OpenStudy (animelover98):

Yes

OpenStudy (anonymous):

Perhaps you are meant to use another method, but quadratic formula is easiest.

OpenStudy (animelover98):

Alright, I understand the quadratic formula. How would I set it up? Thanks!

OpenStudy (anonymous):

\[ x^2 + 2x + 8 = (1)x^2+(2)x+(8) \]

OpenStudy (anonymous):

\[ a=1,b=2,c=8 \]

OpenStudy (anonymous):

\[ x = \frac{-b\pm \sqrt{b^2-4ac}}{2a} = \frac{-(2)\pm \sqrt{(2)^2-4(a)(8)}}{2(1)} \]

OpenStudy (animelover98):

I ended with \[-2\pm \sqrt{-14}\]

OpenStudy (phi):

did you divide by 2 ? (2a)

OpenStudy (anonymous):

That is correct, but you can simplify it.

OpenStudy (animelover98):

Oh! I see!

OpenStudy (anonymous):

Oh, yeah, wait, divide by 2.

OpenStudy (animelover98):

\[-1\pm \sqrt{-7}\]?

OpenStudy (anonymous):

not quite

OpenStudy (phi):

how and what did you get for inside the square root ? can you show your steps?

OpenStudy (animelover98):

Well, I multiplied 2^2 and got 4. Then I multiplied 4(1)(8) and got 32. I subtracted 4-32.

OpenStudy (animelover98):

I got -28 and divided that by 2 to get -14. @phi

OpenStudy (anonymous):

\[ \frac{\sqrt{-28}}{2}=\frac{\sqrt{-28}}{\sqrt{4}} = \sqrt{-28/4} =\sqrt{-7} \]

OpenStudy (phi):

ok with \( \sqrt{-28} \) but you don't divide the "stuff" inside by 2 however, you can write -28= -4*7 so sqr(-28) = sqr(4*-7) = sqr(4) * sqr(-7)

OpenStudy (phi):

or do as wio showed: to "move" the 1/2 inside the square root, square it to get 1/4 and move it inside to get sqr(-28/4)

OpenStudy (animelover98):

Alright, so to do what you did Wio, I just move it inside the square root and square it. Getting a 4 rather than a 2. Thanks so much, both of you. I think I really understand this better. It took 20 minutes, thank you so much for your time! I can't thank you enough!

OpenStudy (phi):

so you should get \[ \frac{-2 \pm \sqrt{-28}}{2} \\ = -1 \pm \sqrt{-7} \\ = -1 \pm \sqrt{7}\ i \]

OpenStudy (animelover98):

Yes! I did! Thank you both!